OpenStudy (anonymous):

i need help with this question pleaYou own a swimming pool elliptical in shape. You wish to cover it at night to retain its warmth. If the pool measures 18 ft. across at its minor axis and 25 ft. across at its major axis, how much plastic material will you need to cover the pool, assuming you need a 1.5 ft. overhang all around the pool? Area (to the nearest tenth) = _______ sq. ft. se.

OpenStudy (anonymous):

elipse=3.14x A x B formula

OpenStudy (anonymous):

Major axis + 2 * 1.5 = 28 ft Minor axis + 2 * 1.5 = 21 ft Area A = 3.14/2 * 28/2 * 21 = 461.8 ft²

OpenStudy (anonymous):

where did you get 2

OpenStudy (anonymous):

Which " 2" ?

OpenStudy (anonymous):

2x1.5

OpenStudy (anonymous):

Left and right the axis!

OpenStudy (anonymous):

because it has 2 sides right

OpenStudy (anonymous):

" assuming you need a 1.5 ft. overhang all around the pool "

OpenStudy (anonymous):

ok thank you